OpenStudy (anonymous):

solve the equation: 3/(x+1) = 1/(x^2-1)

OpenStudy (mertsj):

\[\frac{3}{x+1}=\frac{1}{x^2-1}\] Cross multiply

OpenStudy (anonymous):

Notice x^2 - 1

OpenStudy (mertsj):

\[3x^2-3=x+1\]

OpenStudy (mertsj):

\[3x^2-x-4=0\]

OpenStudy (mertsj):

\[(3x-4)(x+1)=0\]

OpenStudy (mertsj):

\[3x-4=0\]

OpenStudy (mertsj):

\[x=\frac{4}{3}\]

OpenStudy (mertsj):

\[x+1=0\] \[x=-1\]

OpenStudy (mertsj):

But of course x cannot be -1 because then x^2-1 (the denominator) would be 0 so discard that and x = 4/3

OpenStudy (anonymous):

Thank you. Can you help me with this one? 3/(x) = 12/x+7

OpenStudy (mertsj):

Same thing. Cross multiply. What do you get?

OpenStudy (anonymous):

3x+21=12x

OpenStudy (mertsj):

\[12x=3x+21\]

OpenStudy (mertsj):

Yes. Not finish it up.

OpenStudy (anonymous):

3x-12x+21 -9x=21

OpenStudy (mertsj):

That would be positive 9x

OpenStudy (anonymous):

oh

OpenStudy (mertsj):